Number of people (millions) 14 Category Institutionalized Full-time workers Less than 16 years of age Armed forces (noncivilian) Part-time workers Population Looking for work 113 40 20 27 314 16 Refer to the table above. What is the labor force participation rate for this economy? (Round to the nearest one-tenth of a percent.)
Answer #1

Labor participation rate = labor force/adult population*100

Labor force = employed+unemployed

Employed = 113+27 = 140

Unemployed = 16

Labor force = 140+16 = 156

Adult population = 314-14-40-20 = 240

LFPR = 156/240*100

= 65%